# Evaluating Long Range Dependency Handling in Code Generation Models using Multi-Step Key Retrieval

### v1

@@ Abstract
As language models support larger and larger context sizes, evaluating their ability to make effective use of that context becomes increasingly important. We analyze the ability of several code generation models to handle long range dependencies using a suite of multi-step key retrieval tasks in context windows up to 8k tokens in length. The tasks progressively increase in difficulty and allow more nuanced evaluation of model capabilities than tests like the popular needle-in-the-haystack test. We find that performance degrades significantly (up to 2x) when a function references another function that is defined later in the prompt. We also observe that models that use sliding window attention mechanisms have difficulty handling references further than the size of a single window. We perform simple prompt modifications using call graph information to improve multi-step retrieval performance up to 3x. Our analysis highlights different facets of long-context performance and is suggestive of prompt construction strategies for code completion tools.
@@ 1 Introduction
Long context inference is an increasingly important differentiator of LLMs, with model releases supporting larger and larger context windows [1, 2, 3, 4, 5]. This is enabled by advances in efficient attention implementation such as FlashAttention [6], Grouped-Query Attention [7] and Paged Attention [8], as well as scaling attention by introducing sparsity via windowing [9, 10].
Applications, such as in-editor code completion from tools like GitHub’s Copilot [11] and SourceGraph’s Cody [12], benefit from long context support as they leverage retrieval-augmented generation strategies [13] to incorporate code snippets from across the user’s project into a single prompt. This allows completions to be driven by the user’s context rather than just by what the model learned at training time. Works such as Zhang et al. [14] and Shrivastava et al. [15] have proposed various approaches to improve cross-repository code completion.
These applications rely heavily on the model’s ability to effectively use everything in its context window, and thus evaluation of long context usage has been a topic of interest in the community. Khandelwal et al. [16] showed that early LSTM-based language models only roughly modelled information from more distant tokens. Recently, the “needle-in-the haystack” test, which measures recall for chat-style LLMs [17] has grown popular, and a parallel key-retrieval method has been used by Rozière et al. [18] to measure recall in auto-complete style code generation models. We believe that desirable long context inference also includes the ability to reason over multiple pieces of information in the input, and needle-in-haystack evaluation approaches do not capture a model’s ability to do this. In the paper we contribute:
@@
    A set of multi-step key retrieval tasks that progressively increase in difficulty and allow evaluation of long-range dependency handling and basic reasoning through indirection.  [^1] We open source the code to generate these tasks at https://github.com/apple/ml-key-retrieval-code-tasks.
@@
    Empirical evaluation of several open source code generation models that reveal architectural choices that affect ability to handle long-range dependencies. We also discover that the order of function declarations has a large effect on model ability to complete these tasks.
@@
    Preliminary recommendations for improving multi-step recall that rely only on prompt modification using information obtainable via existing non-LLM techniques. These include rewriting code to bring dependencies closer together and adding annotations of function dependencies in the form of comments.
We focus on examining how well long-range dependencies are handled in auto-complete style code generation models. While instruction-tuned/chat-style models are an increasingly popular mode of LLM interaction, we believe that auto-complete models are still relevant for applications that need frequent, low latency generation of code.
@@ 2 Related Work
Symbolic reasoning: Zhang et al. [19] study neural networks’ ability to handle indirection by introducing the Pointer Value Retrieval (PVR) task. They train models to take in a sequence of tokens (typically digits), where the first token acts as pointer to the value to be retrieved. The synthetic nature of this task provides full control over the complexity of the problem. Abnar et al. [20] extend this work to add recursive steps to the PVR task, evaluate it in the context of vision transformers, and present a new architecture that supports adaptive compute.
@@
Multi-hop QA: Min et al. [21] study a popular multi-hop reasoning benchmark (HotpotQA) and find that many of the questions can be answered with single hop reasoning. They find many questions embed contextual info that make the task easier, and that weak distractors may make picking the right answer ‘obvious’ without need for the desired reasoning steps. Chen and Durrett [22] further explore dataset design choices to induce multi-hop reasoning.
@@
Long context retrieval: Liu et al. [23] Examine instruction-tuned LLM performance for synthetic key-retrieval tasks as well as multi-document QA and describe a “lost-in-the-middle” phenomenon where information becomes harder to retrieve if it is not near the beginning or end of the prompt.
@@
Long context dependencies: Yu et al. [24] present a code completion benchmark named CoderEval, where problem solutions have varying levels of dependencies on code in the surrounding context. They find that model performance for functions with dependencies is significantly worse than performance for standalone functions. In the context of natural-language Kuratov et al. [25], Yuan et al. [26] and Levy et al. [27] develop benchmarks to evaluate how well language models extract and reason over distributed facts in the presence of noise text. Hsieh et al. [28] develop a rich set of tasks to evaluate long context performance with different dependency structures in natural language.
@@
Chain-of-thought prompting: Wei et al. [29] and Kojima et al. [30] explore prompting strategies to induce multi-step reasoning in instruction tuned models. They find that models perform better on various tasks when prompted to output intermediate steps in the reasoning chain.
@@
While most of the existing literature focuses on chat-style natural language generation, our work focuses on non-instruction-tuned models in the context of autocomplete-style code generation, which are not able to make use of ‘scratchpad’ methods such as chain-of-thought to improve their reasoning. We situate our work in between the PVR work of Zhang et al. [19], which is highly controllable but uses a fully synthetic task, and the code-with-dependencies benchmarking work exemplified by Yu et al. [24], which is more realistic but also makes it more difficult to run controlled experiments that allow discovery of specific failure modes.
@@ 3.1 Task Design
To study long-range dependency handling we propose four multi-step key retrieval tasks of increasing complexity (one-step, two-step, three-step, and concatenation retrieval). These extend the key-retrieval task in Rozière et al. [18] and test models’ ability to integrate multiple pieces of information spread throughout a long context window to make a completion.
@@
Figure 1: Key retrieval tasks with increasing levels of difficulty. Function names and return values are randomized in the actual prompt. See Appendix F for a complete example.
One-step retrieval is equivalent to the key-retrieval task in [18]. We differ from their design by using random strings to construct function names rather than fixed function names, and return values that are string literals rather than integer literals. Using a simple template system, we first generate a key function that returns a random string. The model is then asked to complete an assert statement on the return value of this function. All return strings that are $10$ characters long and all function names are between $13$ and $20$ characters long.
@@ 3.2 Avoiding Reliance on Parametric Knowledge and Trivial Solutions
In our early experiments we observed that models may produce trivial solutions such as assert foo() == foo(). While these are technically correct , such solutions prevent us from testing the models’ ability to retrieve information from the context. To prevent this, we guide the decoding of tokens to make sure that the response starts as a string literal. As we sample output tokens, we use the prefix_allowed_tokens functionality in the HuggingFace transformers library [31] to ensure that the output starts with any number of spaces followed by a single or a double quote. Once an initial quote mark is produced, generation proceeds unrestricted.
@@ 3.3 Long Context Construction
To turn the tasks described above into long context inference problems we add irrelevant snippets to the context window from two sources. Motivated by findings in Min et al. [21] and Chen and Durrett [22] that show that the presence of good distractor functions are important when measuring multi-step reasoning, we first generate a number of synthetic distractor functions that use the same template as the the key and value functions that contain the task information. Then we sample standalone Python functions from the HumanEval dataset [32] to fill out the context window to our desired size. Algorithm 1 describes this dataset generation process in more detail.
@@
In our experiments, we set $n_{k}$ to $100$ for one-step task, and $20$ for other tasks; $n_{d}$ to 00, $1$, and $5$ for three distractor conditions; $n_{p}$ to $\infty$, $150$, $50$, $50$ for one-step, two-step, three-step, and concatenation tasks respectively (this results in a maximum of 6000 prompts for each condition). We set $n_{t}$ to 2k, 4k and 8k tokens.
@@ 3.4 Models
We evaluate these tasks on five open source models: StarCoderBase-1B, StarCoderBase-7B, StarCoderBase-15.5B [33], StarCoder2-7B [34], and Mistral-7B-v0.1 [4]. The StarCoder family of models are competitive on code generation benchmarks and available in a number of different parameter counts. Mistral is a general purpose LLM that is nonetheless competitive on code generation benchmarks. We use implementations from the HuggingFace transformers library [31].
@@
Model HumanEval [^2] Mistral-7B was trained with 8192 context size but reports support for up to 131k tokens. Max Context Size33footnotemark: 3 Sliding Window Size
Mistral-7B 30.5 8192 / 131072 4096
@@
11footnotetext: HumanEval scores for StarCoder* models are from Lozhkov et al. [34], StarcoderBase-1B scores are from https://huggingface.co/bigcode/starcoderbase-1b. Mistral-7B scores are from Jiang et al. [4].
@@ 3.5.1 Overall Task Performance
Our primary metric is accuracy@k, which is similar to the pass@k metric introduced in Chen et al. [32], the only difference is instead of running unit tests, we simply check if the string literal produced is the expected string. We compute accuracy@3 over 10 generations for each input prompt. Hyperparameters for generation are in Appendix C
@@
Task difficulty: 2(a) shows that tasks increase in difficulty in the following order: one-step, two-step, three-step, and concatenation. We note particularly weak performance on the concatenation task. Another factor affecting difficulty is the number of distractor functions. For the more difficult tasks we see a with a smooth degradation in performance as distractors are added (Figure 3). We focus the rest of our analysis on the five distractor condition as it allows us to see a good variation in task performance.
Context size: In 2(b) we observe a small performance drop as context sizes increases for StarCoderBase-1B, StarCoderBase-7B, and StarCoderBase-15. For StarCoder2-7B and Mistral-8B, which both use sliding window attention, a large drop in performance occurs when moving from 4k to 8k context length, we hypothesise that this is related to the size of the sliding window used and will explore this in more detail in Section 3.5.2.
Incorrect responses: We did some manual inspection of the incorrect generations from models and note a number of distinct failure modes. Firstly a noticeable amount of incorrect generations are return values of distractor functions. Secondly we see completely incorrect answers, such as "$3$", or "text". The third failure mode we notice is partially correct results, particularly for the concatenation task where models often fail to produce the second half of the expected string.
@@ 3.5.2 Effect of Task Snippet Position
Figure 4: Effect of key function position on one-step task with 5 distractors. Position is defined as index of the first token in the function normalized by the total number of tokens in the prompt, and grouped into 20 bins. 0 is the beginning of the prompt and 1 is at the end.
In the one-step retrieval task we see good performance with respect to key function position across the entire context window with two notable exceptions. At the 8k context size Mistral-7B-v0.1 and StarCoder2-7B are unable to perform the one-step task when the key function is more than ~4k tokens away from the generation site Figure 4. Both models use a sliding window attention mechanism with a 4k window size. While this mechanism theoretically allows the model to scale very large context sizes (Jiang et al. [4] reports a theoretical attention span of 131k tokens), our results indicate that the model fails to retrieve precise information at a distance greater than the sliding window.
Figure 5: Effect of key and value function position for two-step tasks. StarCoderBase-7B (top), StarCoder2-7B (bottom) with 5 distractors. Color represents accuracy@3 score.
The two-step task has two relevant task snippet positions, Figure 5 shows heatmaps of task performance vs. both key function position and value function position. We see that performance is worse when the key function appears before the value function in the prompt. We call this a forward reference and it turns out this has a large effect on performance. We also observe that performance is generally higher along the diagonal, i.e., when the functions are closer to each other.
For three-step and concatenation tasks, it becomes harder to directly analyze position effects, as there are three task-relevant code snippets with independent positions. We instead explore performance by forward references and spread between task-relevant snippets in the following sections.

### v2

# Evaluating Long Range Dependency Handling in Code Generation LLMs

@@ Abstract
As language models support larger and larger context sizes, evaluating their ability to make effective use of that context becomes increasingly important. We analyze the ability of several code generation models to handle long range dependencies using a suite of multi-step key retrieval tasks in context windows up to 8k tokens in length. The tasks progressively increase in difficulty and allow more nuanced evaluation of model capabilities than tests like the popular needle-in-the-haystack test. We find that performance degrades significantly for many models (up to 2x) when a function references another function that is defined later in the prompt. We also observe that models that use sliding window attention mechanisms have difficulty handling references further than the size of a single window. We perform simple prompt modifications using call graph information to improve multi-step retrieval performance up to 3x. Our analysis highlights ways that long-context performance needs deeper consideration beyond retrieval of single facts within a document.
@@ 1 Introduction
Long context inference is an increasingly important differentiator of LLMs, with model releases supporting larger and larger context windows (Anthropic, 2023; OpenAI, 2023; Google, 2024; Jiang et al., 2023; Touvron et al., 2023). This is enabled by advances in efficient attention implementation such as FlashAttention (Dao et al., 2022), Grouped-Query Attention (Ainslie et al., 2023) and Paged Attention (Kwon et al., 2023), as well as scaling attention by introducing sparsity via windowing (Child et al., 2019; Beltagy et al., 2020).
Applications, such as in-editor code completion from tools like GitHub’s Copilot (Choi, 2024) and SourceGraph’s Cody (Isken & Hill, 2024), benefit from long context support as they leverage retrieval-augmented generation strategies (Gao et al., 2024) to incorporate code snippets from across the user’s project into a single prompt. This allows completions to be driven by the user’s context rather than just by what the model learned at training time. Works such as Zhang et al. (2023) and Shrivastava et al. (2023) have proposed various approaches to improve cross-repository code completion.
These applications rely heavily on the model’s ability to effectively use everything in its context window, and thus evaluation of long context usage has been a topic of interest in the community. Khandelwal et al. (2018) showed that early LSTM-based language models only roughly modelled information from more distant tokens. Recently, the “needle-in-the haystack” test, which measures recall for chat-style LLMs (Kamradt, 2024) has grown popular, and a parallel key-retrieval method has been used by Rozière et al. (2024) to measure recall in code generation models. However, we believe that desirable long context inference also includes the ability to reason over multiple pieces of information in the input, and needle-in-haystack evaluation approaches do not capture a model’s ability to do this. In the paper we contribute:
@@
    A set of multi-step key retrieval tasks that progressively increase in difficulty and allow evaluation of long-range dependency handling and basic reasoning through function calls.  [^1] We open source the code to generate these tasks at https://github.com/apple/ml-key-retrieval-code-tasks.
@@
    Empirical evaluation of several open source and proprietary code generation models. We find that model performance varies greatly depending on the number of steps involved, and the distinctiveness of the target fact compared to the rest of the context. We also discover that the order of function declarations has a large effect on model ability to complete these tasks. We further observe that sliding window mechanisms degrade models’ ability to resolve references beyond the size of the window.
@@
    An investigation of methods for improving multi-step recall that rely only on prompt modification using information obtainable via existing non-LLM techniques, in particular we add annotations of function dependencies in the form of comments.
@@ 2 Related Work
Symbolic reasoning: Zhang et al. (2022) study neural networks’ ability to handle indirection by introducing the Pointer Value Retrieval (PVR) task. They train models to take in a sequence of tokens (typically digits), where the first token acts as pointer to the value to be retrieved. The synthetic nature of this task provides full control over the complexity of the problem. Abnar et al. (2023) extend this work to add recursive steps to the PVR task, evaluate it in the context of vision transformers, and present a new architecture that supports adaptive compute.
@@
Multi-hop QA: Min et al. (2019) study a popular multi-hop reasoning benchmark (HotpotQA) and find that many of the questions can be answered with single hop reasoning. They find many questions embed contextual info that make the task easier, and that weak distractors may make picking the right answer ‘obvious’ without need for the desired reasoning steps. Chen & Durrett (2019) further explore dataset design choices to induce multi-hop reasoning.
@@
Long context retrieval: Liu et al. (2024) examine instruction-tuned LLM performance for synthetic key-retrieval tasks as well as multi-document QA and describe a “lost-in-the-middle” phenomenon where information becomes harder to retrieve if it is not near the beginning or end of the prompt.
@@
Long context dependencies: Yu et al. (2024) present a code completion benchmark named CoderEval, where problem solutions have varying levels of dependencies on code in the surrounding context. They find that model performance for functions with dependencies is significantly worse than performance for standalone functions. In the context of natural-language, Kuratov et al. (2024), Yuan et al. (2024) and Levy et al. (2024) develop benchmarks to evaluate how well language models extract and reason over distributed facts in the presence of noise text. Hsieh et al. (2024) develop a rich set of tasks to evaluate long context performance with different dependency structures in natural language.
@@
Chain-of-thought prompting: Wei et al. (2023) and Kojima et al. (2023) explore prompting strategies to induce multi-step reasoning in instruction tuned models. They find that models perform better on various tasks when prompted to output intermediate steps in the reasoning chain.
@@
While most of the existing literature focuses on chat-style natural language generation, our work focuses on autocomplete-style code generation, where latency requirements often constrain the use of ‘scratchpad’ methods such as chain-of-thought to improve reasoning. We situate our work in between the PVR work of Zhang et al. (2022), which is highly controllable but uses a fairly abstract task, and the code-with-dependencies benchmarking work exemplified by Yu et al. (2024), which is more realistic but also makes it more difficult to run controlled experiments that allow discovery of specific failure modes. We design multi-step retrieval tasks that are similarly centered on indirection, but focus on evaluating inference-time limitations of pre-trained code models, and expand the input design space to include the effects of position and ordering, in addition to the number of hops.
@@ 3.1 Task Design
To study long-range dependency handling we propose four multi-step key retrieval tasks of increasing complexity (one-step, two-step, three-step, and concatenation retrieval). These extend the key-retrieval task in Rozière et al. (2024) and test models’ ability to integrate multiple pieces of information spread throughout a long context window to make a completion.
@@
Figure 1: Key retrieval tasks with increasing levels of difficulty. Function names and return values are randomized in the actual prompt. See Appendix H for a complete example.
One-step retrieval is equivalent to the key-retrieval task in Rozière et al. (2024). We differ from their design by using random strings to construct function names rather than fixed function names, and return values that are string literals rather than integer literals. Using a simple template system, we first generate a key function that returns a random string. The model is then asked to complete an assert statement on the return value of this function. All return strings that are $10$ characters long and all function names are between $13$ and $20$ characters long.
@@ 3.2 Avoiding Reliance on Parametric Knowledge and Trivial Solutions
In our early experiments we observed that models may produce trivial solutions such as assert foo() == foo(). While these are technically correct , such solutions prevent us from testing the models’ ability to retrieve information from the context. To prevent this, we guide the decoding of tokens to make sure that the response starts as a string literal [^2] We are not able to do this guided decoding for the hosted GPT4o models, we instead add an opening quote to the prompt, i.e. assert key_func() == ".. As we sample output tokens, we use the prefix_allowed_tokens functionality in the Hugging Face transformers library (Wolf et al., 2020) to ensure that the output starts with any number of spaces followed by a single or a double quote. Once an initial quote mark is produced, generation proceeds unrestricted.
@@ 3.3 Long Context Construction
To turn the tasks described above into long context inference problems we add irrelevant snippets to the context window from two sources. Motivated by findings in Min et al. (2019) and Chen & Durrett (2019) that show that the presence of good distractor functions are important when measuring multi-step reasoning, we first generate a number of synthetic distractor functions that use the same template as the the key and value functions that contain the task information. Then we sample standalone Python functions from the HumanEval dataset (Chen et al., 2021) to fill out the context window to our desired size. Algorithm 1 describes this dataset generation process in more detail.
@@
In our experiments, we set $n_{k}$ to $100$ for one-step task, and $20$ for other tasks; $n_{d}$ to 00, $1$, and $5$; $n_{p}$ to $\infty$, $150$, $50$, $50$ for one-step, two-step, three-step, and concatenation tasks respectively (this results in a maximum of 6000 prompts for each condition). We set $n_{t}$ to 2k, 4k and 8k tokens.
@@ 3.4 Models
We evaluate these tasks on six open source models: StarCoderBase-1B, StarCoderBase-7B, StarCoderBase-15.5B (Li et al., 2023), StarCoder2-7B (Lozhkov et al., 2024), Mistral-7B-v0.1 (Jiang et al., 2023) and DeepSeekCoder-6.7B-base(Guo et al., 2024). We also evaluate on GPT-4o-mini and GPT-4o22footnotetext: Specifically gpt-4o-mini-2024-07-18 and gpt-4o-2024-11-20. (OpenAI, 2024). The StarCoder family of models are competitive on code generation benchmarks and available in a number of different parameter counts. StarCoder2-7B and DeepSeekCoder-6.7B-base models have been pre-trained on “repository level” data with the explicit aim of improving performance in contexts where there are dependencies between code units. Mistral is a general purpose LLM that is nonetheless competitive on code generation benchmarks. Similarly the GPT-4o models display strong performance on code generation tasks even though they are not specific to code. We use implementations from the Hugging Face transformers library (Wolf et al., 2020) for the open source models.
@@
Model HumanEval [^3] Mistral-7B was trained with 8192 context size but reports support for up to 131k tokens. Max Context Size44footnotemark: 4 Sliding Window Size
Mistral-7B 30.5 8192 / 131k 4096
@@
DeepSeekCoder-6.7B-base 49.4 16384 N/A
GPT-4o-mini 87.2 128k N/A
GPT-4o 90.2 128k N/A
22footnotetext: HumanEval scores for StarCoder* models are from Lozhkov et al. (2024), StarcoderBase-1B scores are from https://huggingface.co/bigcode/starcoderbase-1b. Mistral-7B scores are from Jiang et al. (2023). GPT scores are from https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/
@@ 3.5.1 Overall Task Performance
Our primary metric is accuracy@k, which is similar to the pass@k metric introduced in Chen et al. (2021), the only difference is instead of running unit tests, we simply check if the string literal produced is the expected string. We compute accuracy@3 over 10 generations for each input prompt. Hyperparameters for generation are in Appendix E
@@
Task difficulty: 2(a) shows that tasks increase in difficulty in the following order: one-step, two-step, three-step, and concatenation. We note particularly weak performance on the concatenation task. Another factor affecting difficulty is the number of distractor functions. For the more difficult tasks we see a consistent degradation in performance as distractors are added (Figure 3).
Context size: In 2(b) we observe a small performance drop as context sizes increases for StarCoderBase-1B, StarCoderBase-7B, StarCoderBase-15B, DeepSeekCoder-6.7B, GPT-4o-mini and GPT-4o. For StarCoder2-7B and Mistral-8B, which both use sliding window attention, a large drop in performance occurs when moving from 4k to 8k context length, we hypothesize that this is related to the size of the sliding window used and will explore this in more detail in Section 3.5.3.
3.5.2 Incorrect Responses
We analyzed the incorrect responses from models and note a number of distinct failure modes. Firstly a noticeable amount of incorrect generations are return values of distractor functions. Considering the 1 and 5 distractor conditions, approximately 10% of all responses are distractor answers across models (min=1.8%, max=16.2%). Within just the incorrect responses, approximately 20% are distractors (min=10.9%, max=33.4%). Full details are in Table 7 and Table 8 in the appendix.
For the concatenation task a common failure mode is returning a partial answer (i.e. one of the two values to be concatenated). Approximately 33.1% of all incorrect responses for this task are partial answers from one of the two strings value functions. See Table 9 for details.
To get further insight into other incorrect responses we also conduct an edit distance analysis. This allows us to see if models are mostly getting the answer right and maybe failing on just a few tokens. We compute Levenshtein distance between model responses and ground truth answers for the incorrect responses and find the mean distance to be 10.97. We note that the length of the expected string is 11 in all tasks. This indicates that when the model is wrong it is generally completely wrong rather than just incorrectly generating one or two tokens. Details of this analysis can be found in Table 10
@@ 3.5.3 Effect of Task Snippet Position
We focus the rest of our analysis on the five-distractor condition as it allows us to see a good variation in task performance.
Figure 4: Effect of key function position on one-step task with 5 distractors. Position is defined as index of the first token in the function normalized by the total number of tokens in the prompt, and grouped into 20 bins. 0 is the beginning of the prompt and 1 is at the end. Shaded regions represent 95% confidence intervals
Figure 5: Varied sliding window sizes for Mistral-7B-v0.1 on the one-step task with 5 distractors.
In the one-step retrieval task we see good performance with respect to key function position across the entire context window with two notable exceptions. At the 8k context size Mistral-7B-v0.1 and StarCoder2-7B are unable to perform the one-step task when the key function is more than ~4k tokens away from the generation site Figure 4. Both models use a sliding window attention mechanism with a 4k window size. While this mechanism theoretically allows the model to scale very large context sizes (Jiang et al. (2023) reports a theoretical attention span of 131k tokens), our results indicate that the model fails to retrieve precise information at a distance greater than the sliding window.
Mistral-7B-v0.1 allows for changing the sliding window size at run time. To futher confirm the result above, we measured one-step retrieval performance with sliding windows of 2048 and 8192 (in addition to the default 4096) at the 8k context size, and find that the drop in performance relative to key function position is generally consistent with this result. However we note that even at the 8192 sliding window size, performance begins to drops when the key function is approximately 60% of the size of the prompt away from the generation site, it however does not drop to zero as in the 2048 and 4096 case. Figure 5 shows this in detail.
StarCoderBase-7B
StarCoder2-7B
GPT-4o-mini
Figure 6: Effect of key and value function position for two-step task with 5 distractors. Top-to-bottom StarCoderBase-7B, StarCoder2-7B, GPT-4o-mini. Color represents accuracy@3 score.
The two-step task has two relevant task snippet positions, Figure 6 shows heatmaps of task performance vs. both key function position and value function position. We see that performance is worse when the key function appears before the value function in the prompt for the first two models. We call this a forward reference and it turns out this has a large effect on performance in most models. We also observe that performance is generally higher along the diagonal, i.e., when the functions are closer to each other. We elaborate on this further in Section 3.5.4
For three-step and concatenation tasks, it becomes harder to directly analyze position effects, as there are three task-relevant code snippets with independent positions. We instead explore performance by number forward references, and by spread between task-relevant snippets in the following sections.